# Less is More: Improving LLM Alignment via Preference Data Selection

### v2

@@ 4 Experiments
We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare DM with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of DM in enhancing comprehensive dialogue ability in Section 4.3; next, we explore the online setting in Section 4.4. lastly, we perform an ablation study for the offline preference data selection in Section 4.5.
@@ 4.1 Experimental Setup
Table 1: GPT4 win rates vs. chosen answers on three tasks. DPO is implemented on subsets selected using different filtering strategies. The supervised fintuned model of Llama-3.2-3B-base is used for initialization. Each strategy selected 2,000 samples for DPO training across TL;DR, HH, and Ultrafeedback. Bold and underlined numbers indicate the best and runner-up performances, respectively.
Filter Init Rand External Margin Implicit Margin IFD Margin DM Full
Strategy TOP MID BOT TOP MID BOT TOP MID BOT ADD MUL
TL;DR 34.00 46.50 66.25 42.00 22.00 30.75 43.00 19.75 1.75 41.25 55.25 63.75 83.25 36.75
HH 70.25 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 85.25 87.25 92.00
Ultrafeedback 61.49 82.86 91.18 73.29 25.84 89.81 77.14 37.02 72.05 83.60 54.53 91.55 90.68 80.99
@@
We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (Bai et al., 2022), and (3) UltraFeedback (Cui et al., 2024), which comprises quality-scored model responses across diverse prompts from multiple sources. To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (Meng et al., 2024). In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model. The scores are given by the PairRM (Jiang et al., 2023b) reward model. Statistics about these datasets are in Appendix A.
@@
We perform sample selection and DPO training experiments across three model architectures: Llama-3.2-3B (MetaAI, 2024), Llama-3-8B (Dubey et al., 2024), and Mistral-7B (Jiang et al., 2023a), under Base and Instruct setup. For the base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets. For the instruct setup, we directly use Llama-3-8B-Instruct or Mistral-7B-Instruct-v0.2 as the start of DPO training. Concerning external reward model for margin calculation, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) which is the best reward model at this scale according to the RewardBench leadboard.
Evaluation.
Following Rafailov et al. (2024), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately. For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (Li et al., 2023b) as our evaluation benchmark, which consists of 805 diverse questions.  [^1] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (Ye et al., 2023), whereas AlpacaEval 2.0 employs GPT4-1106-preview. As the ground truth oracle is unavailable, we assess model performance through win rates against reference responses (written by humans or LLMs like GPT4). Specifically, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue. We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments. More details are presented in Appendix A.
@@
For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (Xia et al., 2024)), (2) Conditional Perplexity (IFD) (Li et al., 2023a), which utilizes token probability to measure informativeness. We use the difference in IFD scores between chosen and rejected responses for preference data selection. (3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards. For (2) and (3), We segment the data into TOP (highest positive difference), MID (near-zero difference), and BOT (lowest negative difference) subsets. Specifically, previous work (Wu et al., 2024) argues that low-gap preference pairs are more informative (i.e., chosen and rejected answers are similar), here we use the IFD-MID strategy to quantify this scheme and call it Low-Gap. For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by Xiong et al. (2024); Dong et al. (2024) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
Implementation Details.
The SFT training of the base model is carried out for two epochs with a learning rate of $2\times 10^{-5}$. Sample packing (Tunstall et al., 2024) is employed to accelerate the training, and we use a block size of 4096. As for DPO training, we follow Rafailov et al. (2024) and used a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$. The DPO training process runs for two epochs using the AdamW optimizer. The learning rate follows a cosine schedule with a 0.1 warm-up ratio, and the global batch size is set to 32. For implicit reward calculation, we employ a Llama-3.2-3B SFT model and its DPO-tuned model, which is fine-tuned on 2,000 randomly selected samples from the complete dataset. More implementation details are presented in Appendix B.
Table 2: Performance comparison on AlpacaEval 2.0 using DPO-trained models with different 6,000-sample subsets (10% of full set). Both SFT and Instruct variants of Llama-3-8B were evaluated. LC and WR denote length-controlled and raw win rate, respectively. Bold number denotes the best-performing selected subset.
@@
DM-ADD 18.64 18.46 29.75 29.62 20.98 18.29 35.97 35.79
DM-MUL 19.53 19.09 28.30 28.39 21.67 20.01 36.36 36.47
Firstly, we compare DM and baseline strategies on three widely-used preference datasets: TL;DR, HH, and UltraFeedback. Using a Llama-3.2-3B model as our base architecture, we evaluate different selection methods, each sampling 2,000 training examples for DPO training. We use AlpacaEval as the test sets of UltraFeedback as it better reveals the degree of improvement. The results, measured by GPT4 win rates, are presented in Table 1. We summarize the findings below.
  • DM consistently achieves (near) optimal win rates across all three tasks, while all baselines show weak results on at least one task. This demonstrates that our method is more robust to useless or toxic samples in diverse domains. For example, DM-MUL achieves 17% higher win rates to second most effective baselines on TL;DR task, this is because the DM-MUL assigns low selection priority so long as a low margin is observed from one source of reward.
  • More data in DPO training does not always yield better results - using just 2-5% of carefully selected data can surpass the performance achieved with the full dataset. This finding suggests the presence of ineffective samples in these tasks, as evidenced by the performance patterns of X-MID and X-BOT strategies (where X represents implicit margin). The results in Table 1 reveal that external, implicit, and IFD margins sometimes can identify toxic samples, as models trained on such data perform significantly worse than the original SFT model. This phenomenon underscores the critical importance of data filtering in DPO training. highlighting that training data should undergo a thorough quality assessment before being used in the training process.
  • Among all methods, only implicit margin-BOT consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit rewards into the DM strategy design. Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences. As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
4.2.1 Analysis
We analyze DM from two perspectives: the joint distribution of different sources of reward margins and the training dynamics for subsets with different margin properties. We utilize the UltraFeedback datasets for this analysis.
The left and middle panels of Figure LABEL:fig:visual illustrate the sample distribution across external-implicit and implicit-implicit margin dimensions on UltraFeedback. Several key observations emerge: (1) The correlation between implicit and external reward margins is notably weak. In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa. This pattern is consistent across other datasets (see Appendix C.2). This divergence highlights the distinct preference patterns captured by these two reward types. Supporting this observation, Table 1 demonstrates that neither margin type consistently outperforms the other in subset selection: external margins prove more effective for TL;DR, implicit margins show superior performance on HH, while both perform comparably on UltraFeedback. These findings underscore the importance of combining both reward types in an ensemble approach. (2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B), and the robust correlation is observed on other datasets as well (see Appendix C.2). This validates our design of using a separate small model for implicit reward margin calculations.
The right panel of Figure LABEL:fig:visual shows DPO training loss curves across different subset selection strategies. TOP/BOT selections converge quickly to low loss values, while MID selections maintain steady, higher loss levels. Surprisingly, BOT-selected samples learn as quickly as TOP-selected ones, suggesting that ‘bad’ preferences are also easy to grasp for LLM. The pattern is consistent across different datasets and models (see Appendix D.1 for more results concerning loss and margin curves on other datasets). While this observation might explain proposals that use absolute margin values for selection (Muldrew et al., 2024), Table 1 reveals that TOP and BOT samples ultimately produce opposing effects despite similar training dynamics. However, we observe a difference between BOT area samples selected by external and implicit margins, where Im-BOT exhibits training dynamics that are more similar to Im-TOP area samples in terms of loss reduction speed.
4.3 AlpacaEval 2.0 Win Rate Comparison
@@ 4.2.1 Analysis
In this section, we conduct a detailed examination of the chat task to understand how data filtering enhances both DPO training efficiency and models’ versatile conversational abilities compared to GPT-4-turbo, representing a key application area for preference learning. We use both Llama-3-8B (base) and (instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, a benchmark widely adopted in preference learning studies. The results of DM and various baselines are shown in Table 2.
Table 3: Iterative DPO: AlpacaEval 2.0 benchmark results across three iterations of DPO training with UltraFeedback prompts.
@@
DPO-DM (r1) 16.35 13.09 1624 42.20 45.74 2091
DPO-DM (r2) 23.79 24.17 1901 46.40 50.60 2316
DPO-DM (r3) 32.31 33.91 2565 48.49 54.99 2774
DM-ADD and DM-MUL consistently outperform full dataset DPO training. Table 2 shows models trained on DM-selected subsets (using only 10% of data) achieve 2-4% higher raw and LC win rates compared to full dataset training across different models (base and instruct) and datasets (offline and on-policy). These consistent improvements demonstrate DM’s robustness and effectiveness, validating the value of data filtering for DPO training. In contrast, all baseline methods demonstrate inconsistent performance across different datasets and models. We attribute this instability to ambiguous samples whose margins vary significantly across different reward models, suggesting that proper fusion of reward signals is necessary to mitigate such noise. Appendix D.2 presents additional training loss and margin curves, illustrating distinct training dynamics across different subsets, and the trend is similar to that on the 3B model as in Appendix D.1.
On-policy data primarily benefits its corresponding on-policy model. Llama-3-8B-Instruct shows approximately 8% higher win rates when trained on the on-policy Llama-UltraFeedback data compared to the off-policy UltraFeedback data, while Llama-3-8B-base demonstrates notably smaller improvements between these datasets. This phenomenon persists in both full and subset DPO training.
4.4 Online Iterative DPO
We explore the benefits of data selection in iterative DPO training using prompts from UltraFeedback as in Xiong et al. (2024). Our methodology employs DM-MUL to filter 5,000 samples from a pool of 20,000 generated samples per iteration. We evaluate both the base and instruct versions of Llama-3-8B, conducting single-epoch training rounds with hyperparameters $\beta=0.1$ and learning rate $5\times 10^{-7}$. Table 3 presents our experimental findings, from which we draw two key observations:
While Best-of-N and Worst-of-N preference pairs from a strong reward model are informative, data filtering is still important to DPO training. DPO-DM performs better for both base and instruct settings using only 25% of the data. This can be attributed to the presence of numerous ambiguous, low-margin samples in the constructed online preference dataset. Furthermore, standard iterative DPO shows increasing output length across training rounds, but data filtering mitigates this issue. By round three, DPO-DM demonstrates both superior performance and better output length control.
Table 4: Ablation study on model choice (Mistral-7B-Instruct-v-0.2). Each strategy selects a 6,000-sample subset.
Method LC (%) WR (%) Length
Initial 17.11 14.72 1676
Full 18.00 18.77 4189
Random 19.42 16.64 1621
Low-Gap 13.31 14.35 2525
DM-ADD 23.67 18.21 1527
DM-MUL 26.04 20.53 1568

### v3

@@ 4 Experiments
Table 1: GPT-4 judged win rates for Llama-3.2-3B models fine-tuned with DPO on subsets selected by various data selection strategies. For every strategy and benchmark (TL;DR, HH, UltraFeedback (UF)), 2,000 preference samples were selected. Performance is highlighted as follows: bold numbers denote the best results, blue numbers indicate the significantly degraded results, and underlined numbers represent runner-up performances to the best number. P, Z, and N denote the most positive, near-zero, and most negative selection principles.
Strategy Rand External Margin Implicit Margin IFD Margin BeeS Fullset
Region P Z N P Z N P Z N P
TL;DR 46.50 66.25 42.00 22.00 30.75 43.00 19.75 1.75 41.25 55.25 83.25 36.75
HH 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 90.25 92.00
UF 82.86 91.18 73.29 25.84 89.81 77.14 37.02 72.05 83.60 54.53 91.68 80.99
We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare BeeS with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of BeeS in enhancing comprehensive dialogue ability in Section 4.3. Lastly, we perform diverse ablation studies for the BeeS in Section 4.4.
@@ 4.1 Experimental Setup
We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (volske2017tl, ; stiennon2020learning, ) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (bai2022training, ), and (3) UltraFeedback (cui2023ultrafeedback, ), which comprises quality-scored model responses across diverse prompts from multiple sources. To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (meng2024simpo, ). In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model. The scores are given by the PairRM (jiang2023llm, ) reward model. Statistics about these datasets are in Appendix A.
@@
Our experiments are conducted across four model series: Llama-3.2 (llama32, ), Llama-3 (dubey2024llama, ), Mistral-7B-v2 (jiang2023mistral, ), and Qwen-2.5 (yang2024qwen2, ) under Base and Instruct setups. For the Base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental Instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets. For the Instruct setup, we directly use them as the start of DPO training. Regarding the external reward model, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (liu2024skywork, ) that is the best reward model at this scale according to the RewardBench leadboard. As for the implicit reward model, we employ the Llama-3.2-3B Base and its DPO-tuned model (on 2,000 randomly selected samples from the complete dataset) for $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$.
Implementation and Evaluation.
For DPO training, we follow (rafailov2024direct, ) and use a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$. We run each training for two epochs, with a learning rate of $5\times 10^{-7}$, and a $0.1$ warmup ratio. Following (rafailov2024direct, ), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately. For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (li2023alpacaeval, ) as our evaluation benchmark, which consists of 805 diverse questions. [^2] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (ye2023comprehensive, ), whereas AlpacaEval 2.0 employs GPT4-1106-preview As the ground truth oracle is unavailable, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue. We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments. More details are presented in Appendix A.
@@
For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (xia2024less, )), (2) IFD (li2023quantity, ) (i.e., exponential form of the Point-wise Mutual Information), which measures semantic overlap. We use the difference in IFD scores between chosen and rejected responses for preference data selection. (3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards. For (2) and (3), we segment the data into P (most positive pairs), Z (close to zero pairs), and N (most negative pairs) subsets according to margin values. Specifically, previous work (wu2024beta, ) posits that "hard" preference pairs (where chosen and rejected samples are highly similar) are more beneficial for training, and we use the IFD-Z to quantify this scheme and call it Low-Gap. For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by (xiong2024iterative, ; dong2024rlhf, ) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
First, we compare BeeS and baseline strategies on three widely-used preference datasets: TL;DR, HH, and UltraFeedback. Using a Llama-3.2-3B model as our Base architecture, we evaluate different selection methods, each sampling 2,000 training examples for DPO training. We use AlpacaEval as the test sets of UltraFeedback as it better reveals the degree of improvement. The results, measured by GPT4 win rates, are presented in Table 1. We summarize the findings below.
  • Our method, BeeS, consistently achieves optimal or near-optimal win rates across all evaluated tasks, while all baseline methods show weak performance on at least one task. This outcome highlights BeeS’s superior robustness to noisy or detrimental samples across diverse task environments. Further, more data in DPO training does not always yield better results. Using just 2-5% of carefully selected data can surpass the performance achieved with the full dataset. Insights from Table 1 also reveal the existence of toxic samples and potential pitfalls of certain selection strategies. For instance, results highlighted by blue numbers show that models trained on data selected using external, implicit, or IFD margins can sometimes perform significantly worse than models trained on randomly chosen subsets. Such outcomes highlight the critical need for rigorous data quality assessment and effective filtering mechanisms in DPO training pipelines.
  • Among all methods, only implicit margin-N consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit reward margin into the BeeS strategy. Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences. As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
4.3 AlpacaEval 2.0 Win Rate Comparison
Table 2: Performance comparison on AlpacaEval 2.0 using DPO-trained models with different 6,000-sample subsets (10% of full set). Both SFT and Instruct variants of Llama-3-8B were evaluated. LC and WR denote length-controlled and raw win rate, respectively. Bold number denotes the best-performing selected subset. Blue numbers denote results that show little advantage over random.
@@ 4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
BeeS 19.53 19.09 30.03 30.46 21.67 20.01 36.36 36.47
Full 17.32 15.30 28.64 26.54 19.92 16.45 32.31 32.44
@@
In this section, we aim to understand how data filtering influences DPO training efficiency and models’ versatile conversational abilities, representing a key application area for preference learning. We use both Llama-3-8B (Base) and (Instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, and results are shown in Table 2, LABEL:fig:loss-margin-8b and 6.
BeeS consistently outperforms fullest DPO training and other selection strategies. As shown in Table 2, BeeS-selected subsets achieve around 4% higher win rates compared to full dataset training across all four settings. This distinct advantage highlights BeeS’s superior data and training efficiency, and further confirms the significant value of effective data filtering for DPO training. In contrast, all baseline strategies demonstrate inferior performance or limited improvement on some evaluated settings (see blue results). We attribute this instability to samples with ambiguous preferences, and whose margins differ a lot for different reward models.
Different training dynamics of ‘P’/‘N’/‘Z’ subset region. The left panel of Figure LABEL:fig:loss-margin-8b shows DPO training loss curves for subsets selected by various strategies. Notably, training on subsets filtered according to the ‘P’ and ‘N’ criteria results in a rapid decrease in loss. In contrast, the loss curves corresponding to the ‘Z’ criterion tend to stabilize at consistently higher plateaus. Notably, ’N’-selected samples, which are often assumed as "difficult-to-learn" (gao2025principled, ), can actually be learned as rapidly as ’P’-selected samples, suggesting that ‘bad’ preferences are also easy to grasp for LLM. The pattern is consistent across different datasets and models (see Appendix C.1 for more results. While this observation might explain proposals that use absolute margin values for selection (muldrew2024active, ), Table 1 reveals that ‘P’ and ‘N’ samples produce opposing effects despite similar training dynamics.
The middle and right panels of Figure LABEL:fig:loss-margin-8b illustrate that data subsets selected by BeeS exhibit both the most rapid decrease in training loss and the fastest increase in the DPO training margin, i.e., current train-batch average implicit margin. These concurrent observations of accelerated optimization help to explain the superior performance achieved by BeeS.
@@
DPO-BeeS (r1) 16.35 13.09 1624 42.20 45.74 2091
DPO-BeeS (r2) 23.79 24.17 1901 46.40 50.60 2316
DPO-BeeS (r3) 32.31 33.91 2565 48.49 54.99 2774
Figure 6: AlpacaEval 2.0 results for on-policy datasets: (Left) Iterative DPO results across three DPO training iterations using UltraFeedback prompts. (Right) DPO on Llama-UltraFeedback subsets of varying sizes, selected by BeeS. Results of DPO-variants trained on fullset are also compared.
Extension to Iterative DPO. We explore the data efficiency of iterative DPO using prompts from UltraFeedback as in (xiong2024iterative, ). In comparison, 20k prompts are used for on-policy preference pair generation per iteration, and our online version uses BeeS to reserve only 5k samples per iteration. The results are in the left panel of Figure 6.
There is high redundancy in the on-policy data construction manner. Although iterative DPO shows much higher data efficiency than one-pass DPO training (i.e., better results than those in Table 2), data selection is still important for quality control. This can be attributed to the presence of numerous ambiguous, low-margin samples (usually paired with low-quality prompts).
A smaller $\beta$ value in DPO loss correlates with higher data efficiency. While $\beta$ is commonly recognized as a factor controlling the strength of the Kullback-Leibler (KL) divergence, it also significantly influences data efficiency. Specifically, the DPO loss, defined as $\log\sigma(\beta\times m_{im})$, indicates that a reduced $\beta$ allows for effective gradient updates for more preference pairs with wider margins. To investigate this, we conducted DPO training on the Llama-3-Instruct 8B model using its on-policy dataset, Llama-UltraFeedback, with $\beta$ set to $0.01$. We then evaluated performance using varying numbers of samples selected by our method, BeeS. The results, presented in the right panel of Figure 6, demonstrate that relaxing the margin constraint in the DPO loss substantially improves data efficiency (Refer to Appendix C.4 for results on the Base setup). Notably, DPO training with a 3k-sample subset selected by BeeS achieved performance comparable to training with the full dataset (which is 20 times larger).
Furthermore, we compared BeeS data selection with several established DPO variants that modify the original loss function, including SimPO (meng2024simpo, ), $\beta$-DPO (wu2024beta, ), and $\alpha$-DPO (wu2024alpha, ). Our findings indicate that BeeS is unique in its ability to effectively enhance both the win rate and the length-controlled (LC) win rate. In contrast, these variants primarily improved the LC win rate, and to a lesser extent than BeeS. These results underscore the significant potential of data selection and data efficiency considerations to enhance the original DPO training algorithm.